# Trusted Knowledge Extraction for Operations and Maintenance Intelligence

### v1

@@ 1 Introduction
Existing research in trustworthy KE has produced robust capabilities in several Natural Language Processing (NLP) techniques. NLP-based methods include Named Entity Recognition (NER), which identifies named entities in text and classifies them into a set of entity types Sundheim (1995). In this study, we adopt a multi-stage KE workflow, based on the Information Extraction Pipeline (Bratanic (2021)), which consists of four core NLP tasks: Coreference Resolution (CR), which links different expressions referring to the same entity and has been shown to increase accuracy in many NLP tasks (Sukthanker et al. (2020)); Named Entity Linking (NEL), which enriches identified entities by linking them to unique identities in external knowledge bases [^1] NER is often conflated with Named Entity Disambiguation (NED), which is the task of disambiguating an entity from its possible references, without necessarily linking it to an external KB (Al-Moslmi et al. (2020)).; and Relation Extraction (RE), which identifies meaningful relationships between entities. NER is the fourth task, which is often embedded in NEL and RE, since many NEL and RE systems utilize a multi-stage approach using an NER sub-module to extract entities before linking or relating them, respectively. Together, these steps support the construction of knowledge graphs. The diagram in Figure 1 summarizes the KE Workflow.
@@
To address this gap, we introduce the OMIn [^2] Operations and Maintenance Intelligence, a novel benchmark dataset in the operations and maintenance domain. OMIn is based on curated records from the publicly available FAA Accident/Incident datasets, which shares several peculiarities found in maintenance data: prevalence of rare entities, uncommon or incorrect syntax due to shorthand, abbreviations, acronyms, and small record size. We also release gold standard annotations for NER, CR, and NEL based on OMIn. Note that we did not create a gold standard for RE, for reasons discussed in Section 3.3.4.
@@
The diagram in Figure 2 demonstrates the KE workflow performed on a sample from OMIn.
Figure 2: KE Workflow Applied Example. Here, the KE Workflow is applied to the sentence on the bottom from OMIn to generate the graph on the top. Named entities, like aircraft and Pilot, are denoted in blue to signify the NER subtask in NEL and RE. Then, the CR system (purple) recognizes that the aircraft refers to the same entity in different parts of the sentence, ensuring information relating to the aircraft is consolidated around one node. NEL (green) connects recognized entities to their corresponding Wikidata entries, such as aircraft (Q11436) and aircraft pilot (Q2095549). Finally, RE (red) identifies relationships between entities, with the red edges representing Wikidata properties, such as the pilot operating the aircraft or the brake being a part of the aircraft.
@@
The remainder of this paper is structured as follows: Section 2 reviews related work. Section 3 presents the methodology used in this study, starting with the creation of the OMIn dataset, including gold standards for annotation, then KE tools selection and zero-shot implementation, and finally our evaluation methodology. Section 4 presents the results from evaluating the selected tools on the OMIn dataset. Section 5 discusses the performance, trustworthiness, and technological readiness of the surveyed tools for the maintenance domain based on the results in Section 4. We conclude by summarizing key contributions and listing suggestions for future work.
@@ 2.3 Open Data and Collaboration in Technical Domains
Prior research highlights the strengths and limitations of current approaches to KE, particularly in the integration of LLMs with KGs. Although LLMs excel in generalizability, they often falter in domain-specific accuracy. Our work addresses this gap by introducing the OMIn dataset for the aviation and maintenance sectors, allowing a targeted evaluation of KE tools. In line with open data initiatives (Brundage et al. (2021)), we contribute an open-source dataset and detailed evaluations, setting a benchmark for future research and tool development in this domain.
@@ 3 Methodology
To evaluate KE approaches in the maintenance domain, we construct a benchmark dataset and systematically assess tool performance across four key tasks: NER, CR, NEL, and RE. Figure 3 illustrates the end-to-end workflow, which begins with the collection and curation of domain-specific data and proceeds through gold standard development, tools selection and implementation, and evaluation. Each step is detailed in the subsections that follow.
Figure 3: Conceptual Overview of Methodology Used in This Study. To create the OMIn dataset, we proceed through data selection, pre-processing, and Gold Standards (GS’s) development. Then, we select four tools for each stage of the KE workflow. These sixteen resultant tools are then implemented on the OMIn dataset, each creating a set of results in the form of named entities (NER), co-references (CR), linked entities (NEL), or relational triples (RE). These results are then evaluated against their respective GS’s, or in the case of RE, against qualitative standards for knowledge representation in KGs.
@@ 3.3.1 Named Entity Recognition Gold Standard
Table 2: Distribution of Entity Types across NER Gold Standards. Empty cell values indicate that the Entity Type does not belong to the corresponding standard.
@@ 3.3.4 Absence of Relation Extraction Gold Standard
Unlike the other KE tasks considered in this study, we do not provide a gold standard (GS) for Relation Extraction (RE). The primary impediment to establishing such a GS is the heterogeneity of the relations that different tools recognize. Each evaluated tool is trained on a distinct set of relationships which Table 4 highlight showing illustrative training data relationships for each RE model, as well as a subset of the relation sets they employ. Although some conceptual overlap exists, no unified ontology currently aligns these disparate sets into a coherent framework that could serve as the basis for a universal gold standard.
@@ 3.4 Tools Selection
We selected sixteen tools for evaluation, with four dedicated to each KE task. The primary criterion guiding this selection was the availability of off-the-shelf, open-source solutions specifically fine-tuned and benchmarked for Named Entity Recognition, Coreference Resolution, Named Entity Linking, and Relation Extraction. This approach enables us to compare the tools’ expected in-domain performance with their observed zero-shot results on our maintenance dataset, thereby providing a stable and interpretable baseline.
@@ 3.4.2 Coreference Resolution Tools
  • Autoregressive Structured Prediction with Language Models (ASP) is a T5-based model developed by Google Research in 2022 (Liu et al. ). ASP proposes a conditional language model trained over structure-building actions instead of strings. This enables it to capture the structure of the sentence more effectively and build the target structure step by step. It performs NER, CR, and RE. Its CR models include three based on different sizes of flant5 (base, large, and xl), and one based on T0-3B. It achieved an F1 score of 82.3 for Coreference Resolution on CoNLL-2012.
@@ 3.4.3 Named Entity Linking Tools
We chose three Named Entity Linking (NEL) tools that have achieved state-of-the-art results and are compatible with Wikidata. Additionally, we decided to evaluate the spaCy EntityLinker because of spaCy’s wide recognition.
@@ 3.4.4 Relation Extraction Tools
  • DeepStructWang et al. ) introduced a model pretrained on task-agnostic corpora to enhance language models’ ability to generate structured outputs, such as entity-relation triples, from the text. DeepStruct is a unified model designed to generalize across a wide range of datasets and tasks, including relation extraction, named entity recognition, and event extraction, among others. However, some steps need to be followed to perform inference on a dataset that is not included in DeepStruct’s list of pretrained datasets. The dataset schemes related to the relation extraction task used in the pretrained model are from the datasets NYT, CoNLL04, ADE, and ACE2005. We selected the NYT setup to use in our evaluation, with the highest RE F1 score of 84.6. Additionally, both UniRel and DeepStruct use the NYT dataset, which improves consistency between tools.
  • Packed Levitated Marker (PL-Marker) uses markers in the encoding phase to capture the interrelation between span pairs (Ye et al. (2022)). The novelty of PL-Marker’s approach is its neighborhood-oriented span-packing strategy. PL-marker provides NER and RE models trained on the SciERC and ACE-2005 datasets, as well as three other NER models. Their BERT-based models, which we evaluate, achieved a 69.0 F1 score on ACE-2005 and a 53.2 F1 score on SciERC in relation extraction. We also evaluate their Albert-xxl-based model which achieved state-of-the-art on ACE-2005 relation extraction with an F1 score of 73.0.
3.5 Tools Implementation
We followed the guidelines provided in each tool’s documentation to set up and generate output. To reproduce our results, see our step-by-step methods documented in our repository (Mealey et al. (2024)).
@@ 3.5 Tools Implementation
For tools that ship multiple models, we strived to evaluate the models most applicable to our task. For spaCy tools, we selected the small and large models. For RE tools, we select the best-performing model as well as others we believe would be informative to evaluate. All models are trained on English data. Table 6 summarizes the models implemented for each tool that has more than one model available.
@@
DeepStruct required some additional steps to prepare it for inference on unseen datasets. First, the unseen dataset’s schema must be aligned with the one DeepStruct was trained on to ensure compatibility with the model’s recognized entity and relation types. Additionally, configuration files might need adjustments to properly format the input data. Considering the relation extraction task, DeepStruct schemas are based on datasets like NYT, CoNLL-04, ADE, and ACE-2005. We chose the NYT schema for our evaluation, which produced the highest RE F1 score of 84.6. This also ensures consistency and fairness when comparing DeepStruct with UniRel.
3.6 Evaluation Methodology
We quantitatively evaluated NER, CR, and NEL tools by comparing their outputs to our established gold standard. In contrast, due to the lack of a gold standard for RE tools in the FAA dataset, we employed manual evaluation by a domain expert. This evaluation was guided by the qualitative criteria outlined by Hogan et al. (2021), with a focus on syntactic accuracy, semantic accuracy, and consistency. Further details on the evaluation methodology are provided in Section D. Additionally, we report the total number of triples generated across the entire FAA dataset, as well as the percentage of records for which triples were generated.
@@ 3.6.1 Named Entity Recognition Evaluation
We evaluated each NER tool against the un-typed UTFAA as well as the benchmark-annotated GS corresponding to its training data (e.g., CoNLLFAA for models trained on CoNLL-2003).
For UTFAA, we evaluated entity spans and ignored labels. We report an F1 score with strong-matching and one with weak-matching. In strong-matching, a predicted entity is correct only if it exactly matches a gold standard entity. In weak-matching, a predicted entity is counted correctly if it contains any substring of the gold standard entity, or if the gold standard entity contains any substring of the predicted entity.
In the example: “Narrative: The cargo door was latched before takeoff by Mr. Bowen. Runway conditions at Steven’s Village was extreme,” the entities should be “cargo door”, “takeoff”, “Mr. Bowen”, “runway conditions”, and “Steven’s Village”. However, flair returns “Bowen” and “Steven’s Village” as the entities. Therefore, “Bowen” is marked as incorrect in the strong-matching evaluation, but correct in the weak-matching evaluation since it is a substring of “Mr. Bowen.”
For the benchmark-annotated GSs, we follow SemEval (Segura-Bedmar et al. (2013)) in reporting four F1 metrics: Strict, Type, Exact, and Partial. In Strict and Type, an entity must be labeled with the correct type to be correct. In Strict and Exact, an entity must exactly (strong) match the gold standard entity to be correct, while in Type and Partial, an entity may be a partial (weak) match. Note that Exact and Partial are equivalent to the label-agnostic strong and weak matching used for UTFAA, respectively.
We utilize work by Batista (2018) to implement both NER evaluations.
@@ 3.6.2 Coreference Resolution Evaluation
For CR, we report 4 metrics. This is useful for data sets that are limited in size and is appropriate since maintenance data sets have few coreferences. Our FAA CR Gold Standard only has 18 coreferences, so it is important to have a variety of evaluation metrics.
  • MUC (Vilain et al. (1995)) measures measures link-based correctness.
  • B-CUBED (Bagga and Baldwin (1998)) evaluates the accuracy of individual elements in the clusters.
  • LEA (Moosavi and Strube (2016)) considers the similarity between the predicted and the true clusters which considers precision and recall on a cluster level.
  • CEAF (Luo (2005)) focuses on the links and entities, considering the impact of each decision within the cluster.
We also follow CoNLL-2012 in reporting an unweighted average of MUC, B-CUBED, and CEAF (Pradhan et al. (2012)), and we employed a coreference evaluation tool, corefeval, to conduct our analysis777corefeval: https://github.com/tollefj/coreference-eval.
@@ 3.6.3 Named Entity Linking Evaluation
We evaluate NEL tools in two ways: F1 score and Ontology-based Topological (OT) metrics that use features derived from the structure (topology) of the underlying base ontology.
To calculate the F1 score, we follow Shen et al. (2021) and Usbeck et al. (2015). In addition to the F1 score, we used two OT similarity computations over Wikidata available in the KGTK Semantic Similarity system (USC-ISI-I2 (2021)): Jiang Conrath (JC) and class similarity. More details on how the F1 and OT metrics scores are computed are available in C.
QID Description Label Class JC
Q15766923 scientific journal Fuel 0.03 0.06
Q5507117 short-lived Bay Area post-hardcore musical… Fuel 0.00 0.00
Table 7: Example OT Measurements
For each metric, we also implement three evaluation strategies: strong-matching, weak-matching, and flexible. Strong and weak matching follow the definitions described in Section 3.6.1. The type of matching affects which predicted entities are included in the evaluation set. Flexible evaluation evaluates predicted entities that have a strong match with entities in the Flexible NEL GS, as described in Section 3.3.3. More details on Flexible evaluation are included in C.

### v3

@@ 1 Introduction
Existing research in trustworthy KE has produced robust capabilities in several Natural Language Processing (NLP) techniques. NLP-based methods include Named Entity Recognition (NER), which identifies named entities in text and classifies them into a set of entity types Sundheim (1995). In this study, we adopt a multi-stage KE workflow, based on the Information Extraction Pipeline (Bratanic (2021)), which consists of four core NLP tasks: Coreference Resolution (CR), which links different expressions referring to the same entity and has been shown to increase accuracy in many NLP tasks (Sukthanker et al. (2020)); Named Entity Linking (NEL), which enriches identified entities by linking them to unique identities in external knowledge bases [^1] NER is often conflated with Named Entity Disambiguation (NED), which is the task of disambiguating an entity from its possible references, without necessarily linking it to an external KB (Al-Moslmi et al. (2020)).; and Relation Extraction (RE), which identifies meaningful relationships between entities. NER is the fourth task, which is often embedded in NEL and RE, since many NEL and RE systems utilize a multi-stage approach using an NER sub-module to extract entities before linking or relating them, respectively. Together, these steps support the construction of knowledge graphs. The diagram in Figure 1 summarizes the KE Workflow in our paper.
@@
To address this gap, we introduce the Operations and Maintenance Intelligence benchmark, or OMIn, a novel benchmark dataset in the operations and maintenance domain. OMIn is based on curated records from the publicly available FAA Accident/Incident datasets, which shares several peculiarities found in maintenance data: prevalence of rare entities, uncommon or incorrect syntax due to shorthand, abbreviations, acronyms, and small record size. We also release gold standard annotations for NER, CR, and NEL based on OMIn. Note that we did not create a gold standard for RE, for reasons discussed in Section 4.3.4.
@@
The remainder of this paper is structured as follows: Section 2 reviews related work. Section 3 captures the research problem. Section 4 presents the methodology used in this study, starting with the creation of the OMIn dataset, including gold standards for annotation, then KE tools selection and zero-shot implementation, and finally our evaluation methodology. Section 5 presents the results from evaluating the selected tools on the OMIn dataset. Section 6 discusses the performance, trustworthiness, and technological readiness of the surveyed tools for the maintenance domain based on the results in Section 5. We conclude by summarizing key contributions and listing suggestions for future work.
@@ 2.3 Open Data and Collaboration in Technical Domains
3 Research Problem
Organizations in the aviation and manufacturing domains generate vast amounts of unstructured maintenance records rich with operational insights, yet these records are difficult to analyze due to their inconsistent structure, jargon, and shorthand. KE techniques, such as CR, NER, NEL, and RE, can transform this text into structured formats like knowledge graphs, but most tools are trained on open-domain prose and perform poorly on technical, domain-specific text. Furthermore, prior research highlights the strengths and limitations of current approaches to KE, particularly in the integration of LLMs with KGs. Although LLMs excel in generalizability, they often falter in domain-specific accuracy. This gap presents a major barrier to deploying trusted KE systems in safety-critical settings, where accuracy, reproducibility, and transparency are essential. Progress is further hindered by the lack of public, domain-specific benchmarks; most datasets are proprietary or too narrow for robust evaluation. This work addresses that gap in the aviation and maintenance sectors by introducing the Operations and Maintenance Intelligence benchmark, or OMIn, a benchmark dataset based on FAA incident reports, enabling targeted evaluation of KE tools. In line with open data initiatives (Brundage et al. (2021)), we contribute an open-source dataset and detailed evaluations for zero-shot performance of sixteen KE tools, setting a benchmark for future research and tool development in this domain.
@@ 4 Methodology
To evaluate KE approaches in the maintenance domain, we construct a benchmark dataset and systematically assess tool performance across four key tasks: NER, CR, NEL, and RE. Figure 2 illustrates the end-to-end workflow, which begins with the collection and curation of domain-specific data and proceeds through gold standard development and tools selection, and on to implementation and evaluation. Each step is detailed in the subsections that follow or Section 5, Evaluation, as indicated in the diagram.
Figure 2: Conceptual Overview of Methodology and Evaluation Strategy Used in This Study. To create the OMIn dataset, we proceed through data selection (Subsection 4.1), pre-processing (Subsection 4.2), and Gold Standards (GSs) development (Subsection 4.3). Then, we select four tools for each stage of the KE workflow (Subsection 4.4) These sixteen resultant tools are then implemented on the OMIn dataset, each creating a set of results in the form of named entities (NER), co-references (CR), linked entities (NEL), or relational triples (RE). The experimental setup for this implementation is captured in Subsection 5.1. These results are then evaluated against their respective GSs, or in the case of RE, against qualitative standards for knowledge representation in KGs. Development and implementation of evaluation metrics is discussed in Subsection 5.2.
The outputs of each KE task, entities, co-references, links, and relational triples, capture complementary aspects of knowledge within each OMIn record. Figure  3 illustrates how the extracted results interact with the original data.
Figure 3: Sample KE Tasks Implementation. Here, the KE Tasks are applied to the sentence on the bottom from OMIn to generate the entities and coreferences annotated on the sentence itself as well as the relational triples and links represented in the graph on the top. Named entities, like aircraft and Pilot, are denoted in blue, and may be understood as either direct results from an NER implementation or intermediate results from an NER subtask via NEL and RE. Then, the CR system (purple) recognizes that the aircraft refers to the same entity in different parts of the sentence, ensuring information relating to the aircraft is consolidated around one node. NEL (green) connects recognized entities to their corresponding Wikidata entries, such as aircraft (Q11436) and aircraft pilot (Q2095549). Finally, RE (red) identifies relationships between entities, with the red edges representing Wikidata properties, such as the pilot operating the aircraft or the brake being a part of the aircraft.
@@ 4.3.1 Named Entity Recognition Gold Standard
Table 2: Distribution of Entity Types across NER Gold Standards. The “–” symbol indicates that the Entity Type does not belong to the corresponding standard.
@@ 4.3.4 Absence of Relation Extraction Gold Standard
Unlike the other KE tasks considered in this study, we do not provide a gold standard (GS) for RE. The primary impediment to establishing such a GS is the heterogeneity of the relations that different tools recognize. Each evaluated tool is trained on a distinct set of relationships which Table 4 highlight showing illustrative training data relationships for each RE model, as well as a subset of the relation sets they employ. Although some conceptual overlap exists, no unified ontology currently aligns these disparate sets into a coherent framework that could serve as the basis for a universal gold standard.
@@ 4.4 Tools Selection
We selected sixteen tools for evaluation, with four dedicated to each KE task. The primary criterion guiding this selection was the availability of off-the-shelf, open-source solutions specifically fine-tuned and benchmarked for NER, CR, NEL, and RE. This approach enables us to compare the tools’ expected in-domain performance with their observed zero-shot results on our maintenance dataset, thereby providing a stable and interpretable baseline.
@@ 4.4.2 Coreference Resolution Tools
  • Autoregressive Structured Prediction with Language Models (ASP) is a T5-based model developed by Google Research in 2022 (Liu et al. ). ASP proposes a conditional language model trained over structure-building actions instead of strings. This enables it to capture the structure of the sentence more effectively and build the target structure step by step. It performs NER, CR, and RE. Its CR models include three based on different sizes of flant5 (base, large, and xl), and one based on T0-3B. It achieved an F1 score of 82.3 for CR n on CoNLL-2012.
@@ 4.4.3 Named Entity Linking Tools
We chose three NEL tools that have achieved state-of-the-art results and are compatible with Wikidata. Additionally, we decided to evaluate the spaCy EntityLinker because of spaCy’s wide recognition.
@@ 4.4.4 Relation Extraction Tools
  • DeepStructWang et al. ) introduced a model pretrained on task-agnostic corpora to enhance language models’ ability to generate structured outputs, such as entity-relation triples, from the text. DeepStruct is a unified model designed to generalize across a wide range of datasets and tasks, including RE, NER, and event extraction, among others. However, some steps need to be followed to perform inference on a dataset that is not included in DeepStruct’s list of pretrained datasets. The dataset schemes related to the RE task used in the pretrained model are from the datasets NYT, CoNLL04, ADE, and ACE2005. We selected the NYT setup to use in our evaluation, with the highest RE F1 score of 84.6. Additionally, both UniRel and DeepStruct use the NYT dataset, which improves consistency between tools.
  • Packed Levitated Marker (PL-Marker) uses markers in the encoding phase to capture the interrelation between span pairs (Ye et al. (2022)). The novelty of PL-Marker’s approach is its neighborhood-oriented span-packing strategy. PL-marker provides NER and RE models trained on the SciERC and ACE-2005 datasets, as well as three other NER models. Their BERT-based models, which we evaluate, achieved a 69.0 F1 score on ACE-2005 and a 53.2 F1 score on SciERC in RE. We also evaluate their Albert-xxl-based model which achieved state-of-the-art on ACE-2005 RE with an F1 score of 73.0.
5 Evaluation
5.1 Experimental Setup
We followed the guidelines provided in each tool’s documentation to set up and generate output. To reproduce our results, see our step-by-step methods documented in our repository Mealey et al. (2024a).
@@ 5.1 Experimental Setup
For tools that ship multiple models, we strove to evaluate the models most applicable to our task. For spaCy tools, we selected the small and large models. For RE tools, we select the best-performing model as well as others we believe would be informative to evaluate. All models are trained on English data. Table 6 summarizes the models implemented for each tool that has more than one model available.
@@
DeepStruct required some additional steps to prepare it for inference on unseen datasets. First, the unseen dataset’s schema must be aligned with the one DeepStruct was trained on to ensure compatibility with the model’s recognized entity and relation types. Additionally, configuration files might need adjustments to properly format the input data. Considering the RE task, DeepStruct schemas are based on datasets like NYT, CoNLL-04, ADE, and ACE-2005. We chose the NYT schema for our evaluation, which produced the highest RE F1 score of 84.6. This also ensures consistency and fairness when comparing DeepStruct with UniRel.
5.2 Evaluation Metrics
We quantitatively evaluated NER, CR, and NEL tools by comparing their outputs to our established gold standard. In contrast, due to the lack of a gold standard for RE tools in the FAA dataset, we employed manual evaluation by a domain expert.
@@ 5.2.1 Named Entity Recognition Evaluation
We evaluated each NER tool against the un-typed UTFAA and against benchmark-annotated gold standards aligned with its training data (e.g., CoNLLFAA for models trained on CoNLL-2003). The key difference is that benchmark GS evaluation considers both entity types and spans, whereas UTFAA evaluation considers only spans and ignores types.
  • Strong/Weak (UTFAA): F1 is computed on entity spans only. Strong matching requires exact span agreement; weak matching allows any substring overlap. For example, the gold entity Mr. Bowen is scored incorrect under strong matching if a system outputs only Bowen, but correct under weak matching.
  • Strict/Type/Exact/Partial (benchmark GS): Following SemEval Segura-Bedmar et al. (2013), F1 metrics consider both entity spans and types. Strict and Type require correct entity types; Exact and Partial are label-agnostic. Exact/Partial correspond directly to strong/weak matching on UTFAA.
We used the work of Batista (2018) to implement both NER evaluations.
@@ 5.2.2 Coreference Resolution Evaluation
For CR, we report four standard metrics, which is important given the small size of maintenance datasets: our FAA CR Gold Standard contains only 18 coreferences. Using multiple metrics ensures robust evaluation.
  • MUC (Vilain et al. (1995)): measures link-based correctness.
  • B-CUBED (Bagga and Baldwin (1998)): evaluates the accuracy of individual elements within clusters.
  • LEA (Moosavi and Strube (2016)): computes cluster-level precision and recall based on the similarity between predicted and true clusters.
  • CEAF (Luo (2005)): evaluates alignment of entities and links, considering the impact of each decision within clusters.
  • CoNLL-2012 (Pradhan et al. (2012)): unweighted average of MUC, B-CUBED, and CEAF, following CoNLL-2012.
We used corefeval toolkit [^6] https://github.com/tollefj/coreference-eval to implement all CR evaluations.
@@ 5.2.3 Named Entity Linking Evaluation
We evaluated NEL tools using two complementary approaches: standard F1 score and Ontology-based Topological (OT) metrics derived from the structure of the underlying ontology. We also use three matching strategies, applied to both sets of metrics. Full computation details and worked examples, including Jiang–Conrath (JC) and class similarity, are provided in  C.
  • Metrics:
    • F1: standard precision/recall-based F1, following Shen et al. (2021); Usbeck et al. (2015).
    • Ontology-based Topological (OT) metrics: computed over Wikidata using the KGTK Semantic Similarity system USC-ISI-I2 (2021), including Jiang–Conrath (JC) and class similarity scores.
  • Matching strategies:
    • Strong: predicted entities must exactly match a gold entity.
    • Weak: predicted and gold entities share any substring (see Section 5.2.1).
    • Flexible: evaluates predicted entities that strongly match entities in the Flexible NEL GS (see Section 4.3.3).